OpenStudy (anonymous):

The revenue function for a school group selling n bookmarks is given by R(n)=2n, and the total cost function is given by C(n)=144+0.80n. Determine the number of bookmarks that need to be sold for the school group to break even. Plz help me!!

OpenStudy (anonymous):

I think it's 2n=144+0.80n @xGuardians ???

OpenStudy (xguardians):

What is the cost that they have to break even ? 144?

OpenStudy (xguardians):

@plzhelpMe1111

OpenStudy (anonymous):

If so.. 2-0.80 = 1.2 144/1.2 = 120 n=120

OpenStudy (xguardians):

@Shay17 Yep, goodjob! @plzhelpMe1111 Shay has your answer!